One half my age is 10 years more than one-third my age. How old am I?
Answer by TimothyLamb(4379)

x/2 = x/3 + 10
(x/2 = x/3 + 10)6
3x = 2x + 60
x = 60
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php